OpenStudy (anonymous):

how can i simplify 8x^2y^4z^16 divided by 4xyz^4 ?

OpenStudy (campbell_st):

look at the problem this way \[\frac{8}{4} \times \frac{x^2}{x}\times \frac{y^4}{y} \times \frac{z^{16}}{z^4}\] when dividing the index law is subtract the powers \[\frac{x^a}{x^b} = x^{a-b}\] use this to get the solution and simplify \[\frac{8}{4}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (campbell_st):

hope it helps